# Crowd Comparative Reasoning: Unlocking Comprehensive Evaluations for LLM-as-a-Judge

### v1

@@ Abstract
LLM-as-a-Judge, which generates chain-of-thought (CoT) judgments, has become a widely adopted auto-evaluation method. However, its reliability is compromised by the CoT reasoning’s inability to capture comprehensive and deeper details, often leading to incomplete outcomes. Existing methods mainly rely on majority voting or criteria expansion, which is insufficient to address the limitation in CoT. We propose Crowd-based Comparative Evaluation, which introduces additional crowd responses to compare with the candidate responses, thereby exposing deeper and more comprehensive details within the candidate responses. This process effectively guides LLM-as-a-Judge to provide a more detailed CoT judgment. Extensive experiments demonstrate that our approach enhances evaluation reliability, achieving an average accuracy gain of $6.7\%$ across five benchmarks. Moreover, our method produces higher-quality CoTs that facilitate judge distillation and exhibit superior performance in rejection sampling for supervised fine-tuning (SFT), referred to as crowd rejection sampling, thereby enabling more efficient SFT. Our analysis confirms that CoTs generated by ours are more comprehensive and of higher quality, and evaluation accuracy improves as inference scales.
@@ 1 Introduction
CCE achieves a remarkable average improvement of $6.7\%$ across five evaluation benchmarks, including RewardBench, HelpSteer2, MTBench Human, JudgeBench and EvalBias. When applied to judge distillation, we find high-quality long CoT judgments produced by CCE have a higher efficiency for training a smaller judge model, especially enhancing bias robustness. Moreover, we extend CCE naturally to SFT rejection sampling, refer to crowd rejection sampling, where our approach serves as a quality signal to identify training-efficient samples from the response pool. This extension highlights the reliability and practicality of CCE. Finally, analysis experiments confirm that CCE promotes more comprehensive and deeper CoT reasoning and scales inference effectively.
@@ 4.2 Experiment Result
Table 4: Accuracy of LLM-as-a-Judge on pair-wise comparison benchmarks. CCE can consistently enhance the LLM-as-a-Judge’s performance during the test-time inference phase, especially considerably outperforming maj@16 and 16-Criteria; under the same preference pairs data, the model trained with judgments synthesized using CCE achieves more reliable evaluation results.

### v2

@@ Abstract
LLM-as-a-Judge, which generates chain-of-thought (CoT) judgments, has become a widely adopted auto-evaluation method. However, its reliability is compromised by the CoT reasoning’s inability to capture comprehensive and deeper details, often leading to incomplete outcomes. Existing methods mainly rely on majority voting or criteria expansion, which is insufficient to address the limitation in CoT. We propose Crowd-based Comparative Evaluation, which introduces additional crowd responses to compare with the candidate responses, thereby exposing deeper and more comprehensive details within the candidate responses. This process effectively guides LLM-as-a-Judge to provide a more detailed CoT judgment. Extensive experiments demonstrate that our approach enhances evaluation reliability, achieving an average accuracy gain of $6.7\%$ across five benchmarks. Moreover, our method produces higher-quality CoTs that facilitate judge distillation and exhibit superior performance in rejection sampling for supervised fine-tuning (SFT), referred to as crowd rejection sampling, thereby enabling more efficient SFT. Our analysis confirms that CoTs generated by ours are more comprehensive and of higher quality, and evaluation accuracy improves as inference scales. Our code is available at https://github.com/Don-Joey/CCE.git.
@@ 1 Introduction
CCE achieves a remarkable average improvement of $6.7\%$ across five judge benchmarks, including RewardBench, HelpSteer2, MTBench Human, JudgeBench and EvalBias. When applied to judge distillation, we find that the high-quality long CoT judgments generated by CCE enable a smaller judge model to achieve higher accuracy, yielding an average improvement of $4.5\%$-$5.6\%$ (in Qwen 2.5-7B), particularly enhancing bias robustness. Moreover, we extend CCE naturally to SFT rejection sampling, referred to as crowd rejection sampling, where our approach serves as a quality signal to identify training-efficient samples from the response pool. Our enhanced rejection strategy consistently outperforms both random sampling and vanilla rejection sampling on MTBench and AlpacaEval-v2, demonstrating the reliability and practical utility of CCE in LLM alignment. Finally, our analysis confirms that CCE scales inference effectively and produced CoTs consistently yield more key points and capture finer-grained details within responses compared to Vanilla LLM-as-a-Judge, facilitating more comprehensive and deeper CoT reasoning.
@@ 4.2 Experiment Result
Table 4: Accuracy of CCE using different selection strategies on LLM-as-a-Judge benchmarks. Our proposed Outcome-Removal Criticizing-Selection consistently surpasses performances using other selection strategies during the test-time inference phase.